OpenStudy (anonymous):

A car travels 368 miles in 5 hours and 45 minutes. How many miles does it travel per hour?

OpenStudy (anonymous):

The first step is to change 5hrs and 45 mins to just hours. Are you able to do this?

OpenStudy (anonymous):

5.75?

OpenStudy (anonymous):

Perfect! Now we need to find the speed (miles per hour). So v=d/t what do you get for v?

OpenStudy (anonymous):

64

OpenStudy (anonymous):

Okay so we know that the speed is 64 miles/hour and we want to find the miles traveled in 1 hour. What are we going to do to get this answer?

OpenStudy (anonymous):

Do you have any idea?

OpenStudy (anonymous):

devide by 24?

OpenStudy (anonymous):

well since we know speed v and time t. we can use the equation again so v=d/t --->>>>> that means d=v*t what answer do you get? :)

OpenStudy (anonymous):

5.75

OpenStudy (anonymous):

remember v=64 and t=1 so d=v*t

OpenStudy (anonymous):

So d=64*1=64 So the answer is 64 miles! :D